OpenStudy (anonymous):

What are the solutions of 3x2 = 21x? A. x = 3, x = 7 B. x = 3, x = −7 C. x = 0, x = 7 D. x = 0, x = −7

OpenStudy (anonymous):

C

OpenStudy (hoblos):

3x2 = 21x 3x2-21x=0 3x(x-7)=0 so x=0 or x=7

Directrix (directrix):

C. x = 0, x = 7 x^2 = 7x x^2-7x=0 x(x-7) = 0 Zero Product Property x = 0 or x = 7

OpenStudy (anonymous):

c